# The Hessian of tall-skinny networks is easy to invert

### v2

@@ Abstract
We describe an exact algorithm for solving linear systems $Hx=b$ where $H$ is the Hessian of a deep net. The method computes Hessian-inverse-vector products without storing the Hessian or its inverse in time and storage that scale linearly in the number of layers. Compared to the naive approach of first computing the Hessian, then solving the linear system, which takes storage that’s quadratic in the number of parameters and cubically many operations, our Hessian-inverse-vector product method scales roughly like Pearlmutter’s algorithm for computing Hessian-vector products.
@@ 1 Introduction
The Hessian of a deep net is the matrix of second-order mixed partial derivatives of its loss with respect to its parameters. Decades ago, when deep nets had only hundreds or thousands of parameters, the Hessian matrix could be inverted to implement optimizers that converged much faster than gradient descent [9, 1]. But for large modern deep nets, using the Hessian has become less practical: The Hessian of a model with a billion parameters would have a quintillion entries, which is far larger than can be stored, multiplied, factorized, or inverted, even in the largest data centers. A common workaround is to approximate the Hessian as a low-rank matrix [10, 6] or as a diagonal matrix [2, 3, 5]. Such approximations make it easier to apply the inverse of the Hessian to a vector. This article shows how to compute and apply the inverse of the Hessian exactly without storing the Hessian and without storing its inverse. The Hessian-inverse-vector product can be computed in time and storage that scales linearly with the number of layers in the model, and cubically in the number of parameters and activations in each layer.
@@
Pearlmutter [8] showed the product of the Hessian with a fixed vector (the so-called Hessian-vector product) can be computed in time and storage that scales linearly with the number of layers in the network. The scaling behavior of Pearlmutter’s method is much better than the cubic scaling of the naive algorithm that first computes the Hessian matrix and then multiplies it by the vector. The method transforms the original network into a new network whose gradient is the desired Hessian-vector product. To compute the Hessian-vector product, one then just applies backpropagation to the new network. Given a way to compute the Hessian-vector product, one can indirectly compute the Hessian-inverse-vector product via, say Krylov iterations like Conjugate Gradient. However, the quality of the result would then depend on the conditioning of the Hessian, which is notoriously poor for deep nets [4]. Unfortunately, there seems to exist no variant of Pearlmutter’s trick to compute the Hessian-inverse-vector products directly.
@@
This article offers a way to compute Hessian-inverse-vector products using roughly as much computation and storage as Pearlmutter’s algorithm requires to compute the Hessian-vector product. The Hessian of a deep net has a regular structure that makes it possible to compute Hessian-inverse-vector products exactly and efficiently. Regardless of the specific operations in each layer, the Hessian of a layerwise deep net is a second order matrix polynomial that involves the first order and second order mixed derivatives of each layer, and the inverse of a block-bi-diagonal operator that represents the backpropagation algorithm. Exploiting this structure offers a direct way to compute the Hessian-vector product using exactly the same operations as Pearlmutter’s algorithm (see Appendix B). It also leads to a way to compute Hessian-inverse-vector product that does not require forming or storing the full Hessian. Assume an oracle that offers the first order and second order mixed derivatives of each layer. For an $L$-layer deep net where each layer has $p$ parameters and produces $a$ activations, naively storing the Hessian would require $O(L^{2}p^{2})$ memory, and solving a linear system would require $O\left(L^{3}p^{3}\right)$ operations in addition to the oracle queries. In contrast, we will show how to perform these operations using only $O(L\max(a,p)^{3})$ additional computation. The dependence on the parameters on the number of activations and parameters in each layer is still cubic, but the dependence on the number of layers is only linear. This makes operating on the Hessian of tall and skinny networks more efficient than the Hessian of short and fat networks.
@@ 2 Overview
Our objective is to efficiently solve linear systems of the form $Hx=g$, where $H$ is the Hessian of a deep neural network, without forming $H$ explicitly. To do this, we employ the following strategy:
@@
    Write down the gradient of the deep net in matrix form, as a bi-diagonal system of linear equations. Solving this system uses back-substitution, which in this case coincides exactly with the computations carried out by backpropagation.
@@
    Differentiate this matrix form to obtain an expression for the Hessian. This expression involves a second-order polynomial in the inverse of the aforementioned bi-diagonal matrix.
@@
    Lift this polynomial into a higher-dimensional matrix form. After pivoting, this matrix becomes block-tri-diagonal.
@@
    Factorize the resulting block-tri-diagonal into a lower bi-diagonal block matrix, a diagonal block matrix and an upper block-diagonal matrix (the LDU factorization).

### v3

@@ Abstract
We describe an exact algorithm to solve linear systems of the form $Hx=b$ where $H$ is the Hessian of a deep net. The method computes Hessian-inverse-vector products without storing the Hessian or its inverse. It requires time and storage that scale linearly in the number of layers. This is in contrast to the naive approach of first computing the Hessian, then solving the linear system, which takes storage and time that are respectively quadratic and cubic in the number of layers. The Hessian-inverse-vector product method scales roughly like Pearlmutter’s algorithm for computing Hessian-vector products.
@@ 1 Introduction
The Hessian of a deep net is the matrix of second-order mixed partial derivatives of its loss with respect to its parameters. Decades ago, when deep nets had only hundreds or thousands of parameters, the Hessian matrix could be inverted to implement optimizers that converged much faster than gradient descent [11, 1]. But for large modern deep nets, relying on the Hessian has become impractical: The Hessian of a model with a billion parameters would have a quintillion entries, which is far larger than can be stored, multiplied, or inverted even in the largest data centers. A common workaround is to approximate the Hessian as a low-rank matrix [12, 6] or as a diagonal matrix [2, 3, 5]. Such approximations make it easier to apply the inverse of the Hessian to a vector. This article shows how to compute and apply the inverse of the Hessian exactly without storing the Hessian or its inverse. The Hessian-inverse-vector product can be computed in time and storage that scale linearly with the number of layers in the model, and cubically in the number of parameters and activations in each layer.
@@
Pearlmutter [9] showed how to compute the product of the Hessian with a fixed vector (the so-called Hessian-vector product) in time and storage that scale linearly with the number of layers in the network. This is much faster than the cubic scaling of the naive algorithm that first computes the Hessian matrix and then multiplies it by the vector. His method transforms the original network into a new network whose gradient is the desired Hessian-vector product. To compute the Hessian-vector product, one then just applies backpropagation to the new network. Given a way to compute the Hessian-vector product, one can indirectly compute the Hessian-inverse-vector product via, say Krylov iterations like Conjugate Gradient as proposed by Pearlmutter and more recently re-investigated [8, 10]. However, the quality of the result would then depend on the conditioning of the Hessian, which is notoriously poor for deep nets [4]. Unfortunately, there seems to exist no variant of Pearlmutter’s trick to compute the Hessian-inverse-vector products directly.
@@
The proposed Hessian-inverse-vector product algorithm takes advantage of a deep net’s layerwise structure. Regardless of the specific operations in each layer, the Hessian is a second order matrix polynomial that involves the first order and second order mixed derivatives of each layer, and the inverse of a block bi-diagonal operator that represents the backpropagation algorithm. Multiplying a vector by this structure computes the Hessian-vector product using exactly the same operations as Pearlmutter’s algorithm (see Appendix B). It also leads to a way to compute Hessian-inverse-vector product that does not require forming or storing the full Hessian. To formally characterize the storage and running time of the algorithm, assume an oracle that offers the first order and second order mixed derivatives of each layer. For an $L$-layer deep net where each layer has at most $p$ parameters and generates at most $a$ activations, naively storing the Hessian would require $O(L^{2}p^{2})$ memory, and solving a linear system would require $O\left(L^{3}p^{3}\right)$ operations in addition to the oracle queries. In contrast, we will show how to perform these operations using only $O(L\max(a,p)^{2})$ storage and $O(L\max(a,p)^{3})$ computation in addition to the oracle queries. The dependence on the number of activations and parameters in each layer remains cubic, but the dependence on the number of layers is only linear. This makes operating on the Hessian of tall and skinny networks more efficient than the Hessian of short and fat networks.
@@ 2 Overview
Our objective is to efficiently solve linear systems of the form $Hx=b$, where $H$ is the Hessian of a deep neural network, without forming $H$ explicitly. To do this, we employ the following strategy:
@@
    Write down the gradient of the deep net in matrix form, as a block bi-diagonal system of linear equations. Solving this system uses back-substitution, which in this case coincides exactly with the computations carried out by backpropagation.
@@
    Differentiate this matrix form to obtain an expression for the Hessian. This expression involves a second-order polynomial in the inverse of the aforementioned block bi-diagonal matrix.
@@
    Refactor the expression wit hthe help of auxiliarry variables. These lift the polynomial into a higher-dimensional linear form. After pivoting, this linear form becomes block tri-diagonal.
@@
    Factorize the resulting block-tri-diagonal into a lower block bi-diagonal matrix, a block diagonal matrix, and an upper block diagonal matrix using the LDU factorization.